OpenStudy (anonymous):

1) why do objects with negligible air resistance accelerate toward the earth at the same rate? 2) how can newton's 2nd law be used to explain how air resistance affects the acceleration of an object in free fall? 3) how does newton's 2nd law relate to what a weight scale would read as you move up and down on an elevator?

OpenStudy (anonymous):

1) Because without air resistance both would fall at a constant negative velocity of 9.81, without their mass reacting to air resistance.

OpenStudy (ghazi):

well second law states that force= rate of change of momentum now your acceleration depends on force (F=m*a) and F depends upon change in momentum now if there is an air resistance ....it will slow down the speed of object hence momentum (P=m*V) therefore force is reduced and hence acceleration is less...relatively without resistance

OpenStudy (ghazi):

\[F=\frac{ dP }{ dt }=\frac{ d(m*v) }{ dt }= m \frac{ dv }{ dt }\]=m*a now if body is moving up in an elevator total acceleration of body= (g-a) because acceleration is in opposite to that of the direction of gravity..and in moving downwards acceleration= (g+a)

OpenStudy (ghazi):

@unkabogable clear?

OpenStudy (anonymous):

yup...thanks a lot!! :)

OpenStudy (ghazi):

:) YW